OpenStudy (anonymous):

which equation represents the line that passes through the points (6,-3), (-4,-9)?

pooja195 (pooja195):

Start by finding the slope \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\]

OpenStudy (anonymous):

the slope would be 6/ 1

pooja195 (pooja195):

How did you get that?

OpenStudy (anonymous):

i am getting confused

pooja195 (pooja195):

Ok let's do this step by step :-) \[\LARGE \frac{y_2-y_1}{x_2-x_1}=slope\] \[\LARGE \frac{-9--3}{-4-6}=slope\] \[\LARGE \frac{-9+3}{-4-6}=slope\]

OpenStudy (anonymous):

is it wrong?

pooja195 (pooja195):

It is.

OpenStudy (anonymous):

so it is -6/-10

pooja195 (pooja195):

Yes simplfy it

OpenStudy (anonymous):

the negatives cancel out so it is 6/10 and the simplest form is 3/5

pooja195 (pooja195):

Correct :-) Now are we making an equation in slope intercept form or point slope?

OpenStudy (anonymous):

point slope form

pooja195 (pooja195):

Ok so we have the slope and 2 points given \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

so the equation is y+3=3/5(x-6)

pooja195 (pooja195):

Perfect!

OpenStudy (anonymous):

thanks for helping!

pooja195 (pooja195):

yw ^_^

pooja195 (pooja195):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

thanks! bye:)

pooja195 (pooja195):

bye bye :-)